In a class some of the students are planned for a picnic,each student collect the money twice their number .If the total amount is Rs.7200/- then find the each person share?
 

Grade:12th pass

2 Answers

Rrr
5 years ago
Given that each student collect money twice their number .to find each person share total amount is equal to twice their number
7200=2n
n=7200/2
n=3600
Arun
5 years ago
Let there are n students
 
hence
 
2n + 2n +2n +2n ….....n times = 7200
 
 2 n^2 = 7200
 
n^2 = 3600
 
n = 60 rupees
 
hence per student share is 120 rupees